Clerk A can process 50 applications in 4 hours, and clerk B can process 50 applications in 8 hours. How long will it take both clerks working together to process 100 applications?
Answer by nerdybill(7384)

Clerk A can process 50 applications in 4 hours, and clerk B can process 50 applications in 8 hours. How long will it take both clerks working together to process 100 applications?
Let x = time (hours) it takes for both
then
x(50/4 + 50/8) = 100
multiplying both sides by 8:
x(100 + 50) = 800
x(150) = 800
x = 800/150
x = 80/15
x = 5.33 hours
or
x = 5 hours and 20 minutes